# vCache: Verified Semantic Prompt Caching

### v4

@@ Abstract
Semantic caches return cached responses for semantically similar prompts to reduce LLM inference latency and cost. They embed cached prompts and store them alongside their response in a vector database. Embedding similarity metrics assign a numerical score to quantify the similarity between a request and its nearest neighbor prompt from the cache. Existing systems use the same static similarity threshold across all requests to determine whether two prompts can share similar responses. However, we observe that static thresholds do not give formal correctness guarantees, can result in unexpected error rates, and lead to suboptimal cache hit rates. This paper proposes vCache, the first verified semantic cache with user-defined error rate guarantees. It employs an online learning algorithm to estimate an optimal threshold for each cached prompt, enabling reliable cache responses without additional training. Our experiments show that vCache consistently meets the specified error bounds while outperforming state-of-the-art static-threshold and fine-tuned embedding baselines. We release the vCache implementation and three benchmarks to support future research.
@@ 1 Introduction
We evaluate the effectiveness of vCache in terms of correctness guarantees and overall performance. To assess generalizability, we compare vCache across two embedding models, two LLMs, and five datasets. We find that vCache consistently meets the error-rate bounds and outperforms static threshold baselines, even when using fine-tuned embeddings. Specifically, it achieves up to $2\times$ higher cache hit rates and reduces error rates by up to $6\times$. Our main contributions are as follows:
@@
    We propose vCache, the first semantic cache that enforces a user-defined correctness guarantee by bounding the error rate.
@@
    We introduce three benchmarks, derived from five real-world datasets, that capture common semantic cache use cases: classification tasks, search queries, and open-ended prompt distributions.
@@
    We release the vCache implementation [^1] https://github.com/vcache-project/vCache and three benchmarks  [^2] https://huggingface.co/vCache to support future research.
@@ 2 Related Work
Existing semantic caches, such as GPTCache (Bang, 2023) and industry variants (Razi et al., 2024; Dan Lepow, 2025; lit, 2025; Dasgupta et al., 2025; Li et al., 2024), use a global similarity threshold to make cache hit decisions. This assumes a uniform correlation between similarity and correctness across all prompts and embeddings. However, as illustrated in Figure 3, similarity distributions vary widely, making fixed thresholds unreliable. Further details are provided in Appendix D.6.
Semantic Cache Optimization
The threshold dilemma illustrated in Figure 3 can be addressed via two approaches: optimizing the embedding space or learning more effective thresholding strategies.
Embedding Fine-tuning: Zhu et al. (2024) propose a distillation-based method that fine-tunes embeddings for semantic caching, improving alignment between semantically equivalent prompts and their responses. A related challenge arises in image retrieval, where systems must determine whether a nearest neighbor corresponds to the correct target class. Zhang et al. (2023a) address this by introducing the Threshold-Consistent Margin loss, which enforces tighter intra-class cohesion and clearer inter-class separation by selectively penalizing negative pairs. However, they require supervised training, are limited to open-source embedding models, and can fail to generalize to out-of-distribution data at inference time (Hajipour et al., 2024). vCache’s online learning algorithm does not require training, is model-agnostic, and generalizes to out-of-distribution data (see Appendix E.1).
@@
Semantic Cache Guarantees
Even with high-quality embeddings and a presumably carefully tuned threshold, semantic caches remain inherently approximate. Unless a threshold of 1.0 is used (effectively restricting cache hits to exact prompt matches), there is always a risk of returning incorrect responses (Razi et al., 2024; Jim Allen Wallace, 2024). Despite this, existing semantic caching systems rely on fixed thresholds to decide whether to return a cached response (Dasgupta et al., 2025; Li et al., 2024; Dan Lepow, 2025; Razi et al., 2024; lit, 2025; Bang, 2023). As a result, they offer no formal guarantees on accuracy or error rates, making it difficult to justify their reliability in production environments. To address this, we propose vCache, the first semantic caching system that combines competitive performance with a user-defined correctness guarantee.
@@ 3 Overview of Semantic Caching
Algorithm
Given a prompt, say $x$, we first compute the embedding $\mathcal{E}(x)$ and find its nearest neighbor $\mathrm{nn}(x)$. The caching policy $\mathcal{P}$ then determines whether we should use the cached response for this prompt ($\mathrm{exploit}$) or run the LLM inference ($\mathrm{explore}$). In case we decide to exploit, $r(\mathrm{nn}(x))$ is returned. Otherwise, we run $r(x)=\mathrm{LLM}(x)$, compute $s(x)$ and $c(x)$, update the observations $\mathcal{O}(\mathrm{nn}(x))$ and add $x$ to the database $\mathcal{D}$ using,
@@
Policy of Existing Systems ($\mathbf{\mathcal{P}_{gptCache}(s(x))}$)
In existing semantic caching systems, the decision function is implemented as a fixed threshold rule. Given a user-defined threshold $t$, the cache exploits if $s(x)\geq t$ by returning the cached response $r(\mathrm{nn}(x))$. Otherwise, it explores by invoking the model for a response. As discussed in Section 2, this approach lacks formal guarantees and does not adapt to variation in similarity value distributions.
Policy of vCache ($\mathbf{\mathcal{P}_{vCache}(s(x),\mathcal{O}(\mathrm{nn}(x)),\delta})$)
vCache replaces static thresholding with an embedding-specific decision function that respects a user-defined error bound $\delta$. If the function returns $\mathrm{exploit}$, the cache is sufficiently confident and outputs the cached response $r(\mathrm{nn}(x))$ (Algorithm 1, L5). Otherwise, it returns $\mathrm{explore}$ by inferring the LLM. Section 4 provides further details.
Scope of Definitions
All quantities such as $nn(x)$, $s(x)$, and $c(x)$, are evaluated at a fixed (but arbitrary) point in time. Since the analysis is performed online, these definitions apply consistently across time steps. Policy discussion always refers to a specific embedding in the cache, with all parameters and estimates interpreted as conditional on it.
@@
Figure 3: Results from 45k samples in the SemCacheClassification benchmark. Motivates the need for dynamic, embedding-specific thresholds. Top: Similarity values of correct and incorrect explorations exhibit highly overlapping distributions. Bottom: Optimal per-embedding thresholds vary substantially, indicating that no single threshold can suffice across embeddings (see Appendix D.6).
@@ 4.1 User Guarantee
where $s(x)\in[0,1]$ is the similarity of $x$ with its near neighbour. $t\in[0,1]$ is an embedding-specific decision boundary parameter, and $\gamma>0$ is a parameter controlling the steepness of the function. The sigmoid form is well-suited for this task: it induces a smooth and monotonic relationship between similarity and correctness probability and enables efficient maximum-likelihood estimation (MLE) of the threshold $t$ from labeled data. By fitting a continuous likelihood function rather than enforcing a hard threshold, vCache generalizes better from limited observations.
@@ 5 Evaluation
For our experiments, we use two popular embedding models (GteLargeENv1-5 (Zhang et al., 2024) and E5-large-v2 (Wang et al., 2022)), and two LLMs (Llama-3-8B-Instruct (Dubey et al., 2024) and GPT-4o-mini (gpt, 2024)), representing both high-quality proprietary models and efficient open alternatives. We use the HNSW vector database (Malkov & Yashunin, 2018) with cosine similarity, a standard metric for comparing vector embeddings in semantic caching systems (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025). All experiments are conducted on a machine running Ubuntu 24.04.2 LTS, with an Intel Xeon Platinum 8570 CPU and an NVIDIA Blackwell with 192 GB of memory.
Baselines
We use the following Cache-settings in our experiments
@@
Figure 4: vCache meets the user-defined maximum error rate bound $\delta$ with steadily increasing cache hit rates (vCache is learning). GPTCache shows increasing error and hit rates, illustrating the unreliability of static thresholds. Static baselines use fixed thresholds (0.99 top, 0.86 bottom). See Figure 5 for a threshold vs. $\delta$ Pareto comparison.
Datasets
@@
To the best of our knowledge, no realistic open-source benchmark currently exists for evaluating semantic caches. We introduce and open-source three diverse benchmarks designed to reflect common caching scenarios. Appendix E provides the complete dataset and benchmark cards.
@@
  • SemCacheLMArena: A randomly sampled subset of 60,000 queries from the LM-Arena human preference dataset (Chiang et al., 2024), containing open-ended, user-generated prompts. It models chatbot scenarios where user inputs vary semantically but can require similar responses.
@@
  • SemCacheClassification: A benchmark of 45,000 prompts derived from three classification datasets (Saurabh Shahane, 2023; Talmor et al., 2018; Ni et al., 2019). Simulating environments where LLMs are embedded in SQL queries to perform classification tasks (Liu et al., 2024b).
  • SemCacheSearchQueries: A random subset of 150,000 real-world queries from the ORCAS dataset (Craswell et al., 2020), consisting of short, search-style questions. It reflects scenarios where LLMs are used in web browsers to answer recurring user queries (Wang et al., 2024; Liu et al., 2024a), and caching might reduce redundant inference. The results are presented in Appendix D.1.
Metrics
We measure the following metrics. (1) Error Rate: (Lower is better) defined as $FP/n$, where FP is the number of false positives and $n$ is the total number of prompts. (2) Cache hit rate: (Higher is better) defined as $(TP+FP)/n$ where TP and FP are true positives and false positives, respectively. Together, TP and FP measure the total cache hits. We also show ROC curves Hoo et al. (2017). For the non-deterministic evaluation of vCache (Algorithm 2, Line 5), we compute 95% confidence intervals using Wallis binomial confidence bounds and contingency tests Wallis (2013).
Figure 5: Pareto comparison across a range of thresholds and $\delta$ values. Each point represents a full run on 60k (SemCacheLmArena) or 45k (SemCacheClassification) samples. vCache generally outperforms static-threshold baselines. While it may slightly underperform for small $\delta$, its stability under increasing sample sizes (Figure 4) indicates greater long-term reliability than static thresholds.
vCache Respects User-defined Error-rate Requirements
@@
We evaluate whether vCache satisfies the user-defined error rate bound $\delta$ while maintaining competitive performance. As shown in Figure 4 (left), vCache consistently meets the maximum error rate across $\delta$ values, with actual error remaining below the specified bound. The small gap between maximum error rate and observed error stems from the conservative $t^{\prime}$ estimation, which ensures robustness and can be further refined. Notably, as the error rate stabilizes, vCache continues to increase its cache hit rate (Figure 4, right), demonstrating effective learning over time. In contrast, GPTCache baselines exhibit increasing error rates with sample size, reflecting the inherent limitations of fixed thresholds despite improving hit rates.
@@
We evaluate whether semantic caches benefit from dynamic, embedding-specific thresholds over a single static threshold. To this end, we compare vCache with static-threshold baselines by systematically varying threshold values for GPTCache and the maximum error rate ($\delta$) for vCache across a feasible range. Each point in Figure 5 represents a complete evaluation over the benchmark dataset, enabling direct Pareto comparison between vCache and static-threshold configurations. vCache achieves better ROC curves, higher cache-hit rates at a given error rate, and lower average latency. On SemCacheLMArena (Figure 5, top), it achieves up to 6× lower error, 2× higher hit rates, and reduced latency. On SemCacheClassification (Figure 5, bottom), vCache outperforms all baselines for error bounds above  1.5%, where static methods either violate constraints or underutilize the cache. For bounds below 1.5%, vCache is more conservative, reflecting its strategy of prioritizing correctness by increasing exploration under uncertainty. We note that GPTCache error rates have not fully converged and exhibit an upward trend, suggesting the reported results likely overstate GPTCache performance.
Latency, vLLM, and Additional Baselines
@@
We report additional experiments on embedding generation overhead (Appendix D.2), logistic regression latency overhead (Appendix D.3), vCache performance combined with vLLM (Appendix D.4), and extended baseline evaluations (Appendix D.5).
@@ 6 Limitations
There are two main limitations of vCache. First, for responses longer than a few words, string matching is insufficient and vCache uses a standard LLM-as-a-judge (Zheng et al., 2023a) approach to assess response equivalence (Algorithm 1, L8), requiring an additional LLM inference. However, this can be executed asynchronously outside the critical path, so it does not impact latency (see vCache implementation). Moreover, since the output is a single token (e.g., ”yes” or ”no”), the computational overhead is minimal (Leviathan et al., 2023). In SemCacheLMArena and SemCacheSearchQueries, response equivalence is assessed with LLM-as-a-judge, whereas in SemCacheClassification it is determined by string matching. vCache performs reliably under both evaluation regimes. Second, vCache relies on the assumptions of independent and identically distributed (i.i.d.) data and a sigmoid function family to represent the probability of correctness. If these assumptions are violated, the analysis may not hold. Nonetheless, both assumptions are natural and appear to model most use cases well, as supported by our experimental results.
@@ 7 Conclusion
We introduced vCache, a semantic cache that provides correctness guarantees by learning an optimal similarity threshold for each cached embedding online. This approach addresses the limitations of static thresholds and embedding fine-tuning in the semantic caching domain, ensuring that the error rate remains below a user-specified bound. Our experiments demonstrate that vCache consistently satisfies this guarantee while outperforming existing methods. These results suggest that reliable, interpretable caching for LLMs is both practical and deployable.
Acknowledgement
The authors thank Sebastien Gautier-Jean-Jacques Beurnier and Siavash Ameli for their invaluable contributions. They also thank the Sky Computing Lab at UC Berkeley for their support.

### v5

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@
    We propose vCache, the first predictable semantic cache that enforces a user-defined correctness guarantee by bounding the error rate.
@@
    We introduce four benchmarks, derived from five real-world datasets, that capture common semantic cache use cases: classification tasks, search queries, and open-ended prompt distributions.
@@
    We release the vCache implementation [^1] https://github.com/vcache-project/vCache and four benchmarks [^2] https://huggingface.co/vCache to support future research.
@@ 2 Related Work
Existing semantic caches, such as GPTCache (Bang, 2023) and industry variants (Razi et al., 2024; Dan Lepow, 2025; lit, 2025; Dasgupta et al., 2025; Li et al., 2024), use a global similarity threshold to make cache hit decisions. This assumes a uniform correlation between similarity and correctness across all prompts and embeddings. However, as illustrated in Figure 3, similarity distributions vary widely, making fixed thresholds unreliable. Further details are provided in Appendix F.9.
Semantic Cache Optimization. The threshold dilemma illustrated in Figure 3 can be addressed via two approaches: optimizing the embedding space or learning more effective thresholding strategies.
Embedding Fine-tuning: Zhu et al. (2024) propose a distillation-based method that fine-tunes embeddings for semantic caching, improving alignment between semantically equivalent prompts and their responses. A related challenge arises in image retrieval, where systems must determine whether a nearest neighbor corresponds to the correct target class. Zhang et al. (2023a) address this by introducing the Threshold-Consistent Margin loss, which enforces tighter intra-class cohesion and clearer inter-class separation by selectively penalizing negative pairs. However, they require supervised training, are limited to open-source embedding models, and can fail to generalize to out-of-distribution data at inference time (Hajipour et al., 2024). vCache’s online learning algorithm does not require training, is model-agnostic, and generalizes to out-of-distribution data (see Appendix H.1).
@@
Semantic Cache Guarantees. Even with high-quality embeddings and a presumably carefully tuned threshold, semantic caches remain inherently approximate. Unless a threshold of 1.0 is used (effectively restricting cache hits to exact prompt matches), there is always a risk of returning incorrect responses (Razi et al., 2024; Jim Allen Wallace, 2024). Despite this, existing semantic caching systems rely on fixed thresholds to decide whether to return a cached response (Dasgupta et al., 2025; Li et al., 2024; Dan Lepow, 2025; Razi et al., 2024; lit, 2025; Bang, 2023). As a result, they offer no formal guarantees on accuracy or error rates, making it difficult to justify their reliability in production environments. To address this, we propose vCache, the first semantic caching system that combines competitive performance with a user-defined correctness guarantee.
@@ 3 Overview of Semantic Caching
Algorithm. Given a prompt, say $x$, we first compute the embedding $\mathcal{E}(x)$ and find its nearest neighbor $\mathrm{nn}(x)$. The caching policy $\mathcal{P}$ then determines whether we should use the cached response for this prompt ($\mathrm{exploit}$) or run the LLM inference ($\mathrm{explore}$). In case we decide to exploit, $r(\mathrm{nn}(x))$ is returned. Otherwise, we run $r(x)=\mathrm{LLM}(x)$, compute $s(x)$ and $c(x)$, update the observations $\mathcal{O}(\mathrm{nn}(x))$ and add $x$ to the database $\mathcal{D}$ using,
@@
Policy of existing systems ($\mathbf{\mathcal{P}_{gptCache}(s(x))}$). In existing semantic caching systems, the decision function is implemented as a fixed threshold rule. Given a user-defined threshold $t$, the cache exploits if $s(x)\geq t$ by returning the cached response $r(\mathrm{nn}(x))$. Otherwise, it explores by invoking the model for a response. As discussed in Section 2, this approach lacks formal guarantees and does not adapt to variation in similarity value distributions.
Policy of vCache ($\mathbf{\mathcal{P}_{vCache}(s(x),\mathcal{O}(\mathrm{nn}(x)),\delta})$). vCache replaces static thresholding with an embedding-specific decision function that respects a user-defined error bound $\delta$. If the function returns $\mathrm{exploit}$, the cache is sufficiently confident and outputs the cached response $r(\mathrm{nn}(x))$ (Algorithm 1, L5). Otherwise, it returns $\mathrm{explore}$ by inferring the LLM. Section 4 provides further details.
Scope of definitions. All quantities such as $nn(x)$, $s(x)$, and $c(x)$, are evaluated at a fixed (but arbitrary) point in time. Since the analysis is performed online, these definitions apply consistently across time steps. Policy discussion always refers to a specific embedding in the cache, with all parameters and estimates interpreted as conditional on it.
@@
Figure 3: Results from 45k samples in the SemCacheClassification benchmark. Motivates the need for dynamic, embedding-specific thresholds. Top: Similarity values of correct and incorrect explorations exhibit highly overlapping distributions. Bottom: Optimal per-embedding thresholds vary substantially, indicating that no single threshold can suffice across embeddings (see Appendix F.9).
@@ 4.1 User Guarantee
where $s(x)\in[0,1]$ is the similarity of $x$ with its near neighbour. $t\in[0,1]$ is an embedding-specific decision boundary parameter, and $\gamma>0$ is a parameter controlling the steepness of the function. The sigmoid form is well-suited for this task: it induces a smooth and monotonic relationship between similarity and correctness probability and enables efficient maximum-likelihood estimation (MLE) of the threshold $t$ from labeled data (justification in Appendix E). By fitting a continuous likelihood function rather than enforcing a hard threshold, vCache generalizes better from limited observations.
@@ 5 Evaluation
For our experiments, we use three popular embedding models (GteLargeENv1-5 (Zhang et al., 2024), E5-large-v2 (Wang et al., 2022), and OpenAI text-embedding-3-small (ope, 2025)), and two LLMs (Llama-3-8B-Instruct (Dubey et al., 2024) and GPT-4o-mini (gpt, 2024)), representing both high-quality proprietary models and efficient open alternatives. We use the HNSW vector database (Malkov & Yashunin, 2018) with cosine similarity, a standard metric for comparing vector embeddings in semantic caching systems (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025). All experiments are conducted on a machine running Ubuntu 24.04.2 LTS, with an Intel Xeon Platinum 8570 CPU and an NVIDIA Blackwell with 192 GB of memory.
Metrics. We measure the following metrics. (1) Error Rate: (Lower is better) defined as $FP/n$, where FP is the number of false positives and $n$ is the total number of prompts. (2) Cache hit rate: (Higher is better) defined as $(TP+FP)/n$ where TP and FP are true positives and false positives, respectively. Together, TP and FP measure the total cache hits. We also show ROC curves Hoo et al. (2017). For the non-deterministic evaluation of vCache (Algorithm 2, Line 5), we compute 95% confidence intervals using Wallis binomial confidence bounds and contingency tests Wallis (2013).
Figure 4: vCache meets the user-defined maximum error rate bound $\delta$ with steadily increasing cache hit rates (vCache is learning). GPTCache shows increasing error and hit rates, illustrating the unreliability of static thresholds. Static baselines use fixed thresholds. See Figure 5 for a threshold vs. $\delta$ Pareto comparison.
Figure 5: Pareto comparison across a range of thresholds and $\delta$ values. Each point represents a full run on all samples of a dataset.
Baselines. We use the following Cache-settings in our experiments
@@
We implement additional baselines for ablations and show the results in Appendix F.8.
@@
Datasets. To the best of our knowledge, no realistic open-source benchmark currently exists for evaluating semantic caches. We introduce and open-source four diverse benchmarks designed to reflect common caching scenarios. Appendix H provides the complete dataset and benchmark cards.
@@
  • SemCacheLMArena: Randomly sampled subset of 60,000 queries from the LM-Arena human preference dataset (Chiang et al., 2024), containing open-ended, and user-generated prompts.
@@
  • SemCacheClassification: Benchmark of 45,000 prompts derived from three classification datasets (Saurabh Shahane, 2023; Talmor et al., 2018; Ni et al., 2019).
  • SemCacheSearchQueries: Random subset of 150,000 web-search queries (Craswell et al., 2020).
  • SemCacheCombo: 27,500-prompt benchmark combining SemCacheSearchQueries and distinct SemCacheLMArena queries to model partial workloads with no cache hits.
@@
vCache respects user-defined error-rate requirements. We evaluate whether vCache satisfies the user-defined error rate bound $\delta$ while maintaining competitive performance. As shown in Figure 4, vCache consistently meets the maximum error rate across $\delta$ values, with actual error remaining below the specified bound. The small gap between maximum error rate and observed error stems from the conservative $t^{\prime}$ estimation, which ensures robustness and can be further refined. Notably, as the error rate stabilizes, vCache continues to increase its cache hit rate, demonstrating effective learning over time. In contrast, GPTCache baselines exhibit increasing error rates with sample size, reflecting the inherent limitations of fixed thresholds despite improving hit rates.
@@
Dynamic and embedding-specific thresholds are superior to static thresholds. We evaluate whether semantic caches benefit from dynamic, embedding-specific thresholds over a single static threshold. To this end, we compare vCache with static-threshold baselines by systematically varying threshold values for GPTCache and the maximum error rate ($\delta$) for vCache across a feasible range. Each point in Figure 5 represents a complete evaluation over the benchmark dataset, enabling direct Pareto comparison between vCache and static-threshold configurations. vCache achieves better ROC curves, higher cache-hit rates at a given error rate, and lower average latency. On SemCacheLMArena, it achieves up to 26× lower error, 12.5× higher hit rates, and reduced latency. On SemCacheClassification, vCache outperforms all baselines for error bounds above  1.5%, where static methods either violate constraints or underutilize the cache. For bounds below 1.5%, vCache is more conservative, reflecting its strategy of prioritizing correctness by increasing exploration under uncertainty. We note that GPTCache error rates have not fully converged and exhibit an upward trend (Figure 4), suggesting the reported results likely overstate GPTCache performance.
@@
Convergence, Sigmoid, Latency, vLLM, and Additional Baselines. We report additional experiments on the convergence speed (Appendix D), aggregate empirical correctness probability (Appendix E), $\tau$ computation (Appendix F.4), embedding generation (Appendix F.5), logistic regression (Appendix F.6), vLLM inference (Appendix F.7), and extended baseline evaluations (Appendix F.8).
@@ 6 Limitations
There are two main limitations of vCache. First, for responses longer than a few words, string matching is insufficient and vCache uses LLM-as-a-judge (Zheng et al., 2023a) to assess response equivalence (Algorithm 1, L8), requiring an additional LLM inference. However, this can be executed asynchronously outside the critical path, not impacting latency (see vCache implementation). Since the output is a single token (e.g., ”yes” or ”no”), the overhead is minimal (Leviathan et al., 2023). In SemCacheLMArena and SemCacheSearchQueries, response equivalence is assessed with LLM-as-a-judge, whereas in SemCacheClassification it is determined by string matching. vCache performs reliably under evaluation regimes (Figure 4). Second, vCache relies on i.i.d. data and a sigmoid function family to represent the probability of correctness. If these assumptions are violated, the analysis may not hold. Nonetheless, both assumptions are natural and appear to model most use cases well, as supported by our experimental results (Appendix E).
@@ 7 Conclusion
We introduced vCache, a semantic cache that provides correctness guarantees by learning an optimal similarity threshold for each cached embedding online. This approach addresses the limitations of static thresholds and embedding fine-tuning in the semantic caching domain, ensuring that the error rate remains below a user-specified bound. Our experiments demonstrate that vCache consistently satisfies this guarantee while outperforming existing methods w.r.t error and cache hit rates. These results suggest that reliable, interpretable caching for LLMs is both practical and deployable.
8 Acknowledgment
The authors thank Sebastien Gautier-Jean-Jacques Beurnier and Siavash Ameli for their invaluable contributions. This research is partly supported by gifts from Accenture, Amazon, AMD, Anyscale, Broadcom, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Lightspeed, Mibura, NVIDIA, Samsung SDS, and SAP.
9 Ethics Statement
This work contributes to the advancement of machine learning by improving the efficiency of large language model (LLM) inference. By reducing both computational cost and latency, our approach makes LLM-based systems more accessible to organizations and individuals with limited computational resources, thereby lowering the barrier to adoption. In addition, by decreasing the frequency of full LLM invocations, our method reduces overall compute demand and, consequently, the environmental impact associated with training and operating large-scale AI infrastructure.
10 Reproducibility Statement
We have taken several steps to ensure reproducibility of our work. The full implementation of vCache, including all algorithms and experiments, is publicly available in a GitHub repository [^3] https://github.com/vcache-project/vCache. The four semantic caching benchmarks introduced in this paper are released on HuggingFace [^4] https://huggingface.co/vCache, and Appendix H describes their construction and preprocessing in detail. All experimental settings, including hardware, software, and model configurations, are specified in Section 5. Theoretical assumptions and complete proofs of our correctness guarantees are provided in Appendix C. Together, these resources ensure that our results can be verified and extended by future work.